On a weekend outing, carlos rents a motorboat for 8 hrs to travel down the river and back. the rental operator tells him to go for 3 hrs downstream, leaving 5 hrs to return upstream.
If the river flows at a speed of 7mph how fast must Carlos travel in order to return in 8 hrs?
How far downstream did carlos go before he returned back?

Answer by richwmiller(17219)

r*t=d
(r+7)*3=(r-7)*5,
3r+21=5r-35
56=2r
28 mph=r
3*35=105 miles downstream